A firm develops a new product that will add 1 Million to profit each year for five years. If the discount rate is 10 percent per year, how much will this new product add to shareholder value?
1
Expert's answer
2018-03-08T09:43:07-0500
The Value of a stock is given by Dividend per share divided by discount rate minus growth rate.
Value of stock = (dividend per share)/ (discount rate – growth rate)
Value of stock = {(100,000)/ (5 – 0.1)}* 1/100 =100,000/4.9 = 204.08/100 = 2.0408%